# Approximate independence of permutation mixtures

### v2

@@ Abstract
We prove bounds on statistical distances between high-dimensional exchangeable mixture distributions (which we call permutation mixtures) and their i.i.d. counterparts. Our results are based on a novel method for controlling $\chi^{2}$ divergences between exchangeable mixtures, which is tighter than the existing methods of moments or cumulants. At a technical level, a key innovation in our proofs is a new Maclaurin-type inequality for elementary symmetric polynomials of variables that sum to zero and an upper bound on permanents of doubly-stochastic positive semidefinite matrices. Our results imply a de Finetti-style theorem (in the language of Diaconis and Freedman, 1987) and general asymptotic results for compound decision problems, generalizing and strengthening a result of Hannan and Robbins (1955).
@@ 1 Introduction
  • Under $\mathbb{P}_{n}$, the observation $X$ is of the form $\vartheta+Z$, where $\vartheta$ is uniformly distributed on the subset of balanced vectors in {±$\{\pm\mu\}^{n}$ (i.e., vectors with exactly $n/2$ entries equal to $-\mu$ and $n/2$ equal to $+\mu$) and $Z\sim{\mathcal{N}}(0,I_{n})$ is independent Gaussian noise;
  • Under $\mathbb{Q}_{n}$, the observation $X$ is of the form $\vartheta+Z$, where $\vartheta$ is uniformly distributed on {±$\{\pm\mu\}^{n}$ and $Z\sim{\mathcal{N}}(0,I_{n})$ is independent Gaussian noise. Note that the entries of $X$ under $\mathbb{Q}_{n}$ are themselves i.i.d., and that the marginal distribution of each coordinate of $X$ under $\mathbb{Q}_{n}$ and $\mathbb{P}_{n}$ is the same.
@@
The law of large numbers implies that a random vector $\vartheta$ with i.i.d. {±$\{\pm\mu\}$ entries will be nearly balanced, and, if $\mu\lesssim 1$, the Gaussian noise will make it difficult to detect whether $\vartheta$ is perfectly balanced or only nearly so. It is therefore natural to conjecture that $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ will not be distinguishable from each other as long as $\mu$ is not too large. Surprisingly, however, as we show in Section 2, existing approaches to bounding statistical distances between mixtures fail to establish this fact. The goal of this paper is to develop new strategies for analyzing such high-dimensional mixtures, which achieve tight bounds.
@@
Strikingly, (1) implies that the $\chi^{2}$ divergence between these two high-dimensional models is bounded independent of the dimension $n$—in particular, the sequence $\mathbb{P}_{n}$ is contiguous to $\mathbb{Q}_{n}$ in the sense of Le Cam [LC60], so that asymptotic properties of the mixture $\mathbb{P}_{n}$ can be analyzed under the simpler product measure $\mathbb{Q}_{n}$. The bound in (1) also reveals other features: since $\chi^{2}({\mathcal{N}}(+\mu,1)\|{\mathcal{N}}(-\mu,1))=e^{4\mu^{2}}-1$, the $\chi^{2}$ divergence between the two models is polynomially smaller than the “single-letter” $\chi^{2}$ divergence that would arise from comparing the two univariate marginal distributions that appear in the definition of $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$. As we shall show, up to constants, the bounds in (1) are tight.
@@
The study of permutation mixtures is motivated by important theoretical questions in a number of areas:
Compound decisions and empirical Bayes. In the compound decision problem [Rob51], the statistician observes independent data $X_{1},\dots,X_{n}$, where $X_{i}\sim P_{\theta_{i}}$ for $i=1,\dots,n$; the goal is to estimate the vector of parameters under a separable loss $L(\theta,\widehat{\theta}):=\sum_{i=1}^{n}\ell(\theta_{i},\widehat{\theta}_{i})$. This setting is the starting point of the theory of empirical Bayes [Efr19], which views the parameters as having been drawn i.i.d. from a prior distribution, which is in turn estimated from the observations. An oracle version of this problem was suggested by Brown and Greenshtein [BG09] (see also [JZ09]), in which the statistician knows the unordered set $\{\theta_{1},\dots,\theta_{n}\}$, but not their correspondence to the observations $(X_{1},\dots,X_{n})$. Equivalently, the observations may be viewed as arising from the permutation mixture $\mathbb{P}_{n}$. As Greenshtein and Ritov [GR09] observe, however, most oracle inequalities in the empirical Bayes literature compare to an oracle that knows the unordered set $\{\theta_{1},\dots,\theta_{n}\}$ but is restricted to the use of “simple” decision rules, of the form $\hat{\theta}=(\Delta(X_{1}),\dots,\Delta(X_{n}))$ for some fixed univariate function $\Delta$. Such rules are in fact optimal when the observations arise from not from $\mathbb{P}_{n}$ but from its independent counterpart $\mathbb{Q}_{n}$. The empirical Bayes literature contains several ad hoc techniques for bounding the difference between these two approaches, going back to [HR55]. Our results can be used to give new, general risk bounds in this setting, see Corollary 1.7.
de Finetti-style theorems. The classical de Finetti theorem [DF29] asserts that an infinite exchangeable sequence of binary random variables is a mixture of i.i.d. Bernoulli sequences. Since the work of Diaconis [Dia77], there has been interest in obtaining quantitative forms of this theorem for finite exchangeable sequences [Sta78, DF80, GK21], and, more broadly, to obtain “de Finetti-style theorems” in general settings [DF87]. Our main results, which establish that the exchangeable measure $\mathbb{P}_{n}$ is close to the product measure $\mathbb{Q}_{n}$, are theorems of this type. As Theorem 1.5 shows, our bounds yield new approximation results for general exchangeable sequences.
@@
Information geometry of high-dimensional mixture models. High-dimensional mixtures—such as the measures $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ defined above—are ubiquitous models in theoretical statistics (see, e.g., [LNS99, CL11, IS12, WY16, JHW18]). The analysis of such measures is routine in the univariate case (or, more generally, for product measures), but obtaining sharp bounds is difficult for high-dimensional mixtures whose coordinates are not independent. The setting we study is one in which we can rigorously compare $\mathbb{P}_{n}$ to a simpler, independent counterpart $\mathbb{Q}_{n}$. This comparison may be directly relevant in analyzing mixtures with permutation invariance. Moreover, as we highlight in Section 1.2, our proofs are based on new techniques for bounding divergences between high-dimensional mixtures, which may be of broader applicability.
@@
$\chi^{2}(P\|Q)=\int\frac{(\mathrm{d}P-\mathrm{d}Q)^{2}}{\mathrm{d}Q}\,\qquad H^{2}(P,Q)=\int\left(\sqrt{\mathrm{d}P}-\sqrt{\mathrm{d}Q}\right)^{2}.$
@@
When the $\chi^{2}$ divergence is replaced by the Kullback–Leibler (KL) divergence in Definition 1.1, the $\chi^{2}$ channel capacity coincides with the usual notion of channel capacity in information theory. We have ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})$ by convexity, and $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})\leq{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})+1$ by [PW24, Eqn. (7.33)]. In addition, $H^{2}(P_{1},P_{2})=2$ if and only if $P_{1}$ and $P_{2}$ are mutually singular, in which case it is easy to see that $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ must grow with $n$ in the worst case (cf. Lemma 5.1). Therefore, the quantity $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})$ measures the extent of singularity of ${\mathcal{P}}$. Finally, we note that all quantities (whenever finite) depend only on ${\mathcal{P}}$ but not on the dimension $n$ of $\mathbb{P}_{n}$ or $\mathbb{Q}_{n}$. We present our main theorem below, which upper bounds the $\chi^{2}$ divergence using only the above quantities.
@@
Theorem 1.2 shows that if ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq 1-\delta$ for some constant $\delta>0$, then the $\chi^{2}$ divergence is upper bounded by a dimension-independent constant $O_{\delta}({\mathsf{C}}_{\chi^{2}}({\mathcal{P}})^{2})$. In addition, even if ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\geq 1$ but $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})<\infty$, the $\chi^{2}$ divergence is still upper bounded by a large albeit dimension-independent constant. This shows that the contiguity relation $\{\mathbb{P}_{n}\}\triangleleft\{\mathbb{Q}_{n}\}$ is not unique to the Gaussian family but holds for a large class of probability families ${\mathcal{P}}$. Specializing ${\mathcal{P}}$ to specific classes of distributions leads to the following corollary.
@@
The proof of Corollary 1.3 follows from Theorem 1.2 and the computations of $({\mathsf{C}}_{\chi^{2}}({\mathcal{P}}),\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))$, which we defer to Section C.1. For the example presented at the beginning of the introduction, a weaker upper bound $\exp(O(\mu^{2}))$ of (1) follows from point 2 of Corollary 1.3; we refer to Section 3.1 for the proof of the stronger form (1) (via the inequality ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq 1-\exp(-\mu^{2})$ in (2)). We will also discuss the tightness of these upper bounds in Section 5.1; typically, the quadratic dependence in (4) is tight when ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ is small, and both the base and exponent of (5) are not improvable in general. However, for specific ${\mathcal{P}}$ there could still be gaps, and we refer the discussions to Section 5.1.
Another direct corollary of Theorem 1.2 is the application of the continuity relation $\{\mathbb{P}_{n}\}\triangleleft\{\mathbb{Q}_{n}\}$ to mutual information with permutation structures. Let $(\theta_{1},\dots,\theta_{n})$ be a (deterministic or random) vector, with $P_{\theta_{1}},\dots,P_{\theta_{n}}\in{\mathcal{P}}$ being probability distributions over a space ${\mathcal{X}}$. Let $\vartheta=(\vartheta_{1},\dots,\vartheta_{n})$ be a uniformly random permutation of $(\theta_{1},\dots,\theta_{n})$, i.e. $\vartheta_{i}=\theta_{\pi(i)}$ with $\pi\sim\text{\rm Unif}(S_{n})$, and $X=(X_{1},\dots,X_{n})$ be a random vector with independent coordinates $X_{i}\sim P_{\vartheta_{i}}$ conditioned on $\vartheta$. From the statistical perspective, the random vector $\vartheta$ represents a permutation prior on the model parameter and $I(\vartheta;X)$ denotes the amount of information $X$ provides for estimating $\vartheta$. From the information theoretic perspective, when the vector $\theta$ represents the random message to be communicated, the maximum mutual information $I(\theta;X)$ over all possible distributions of $\theta$ is the channel capacity of the so-called noisy permutation channel [Mak20]. Our contiguity relation implies that the mutual information in both scenarios is close to its counterpart when $\vartheta\sim\left(\frac{1}{n}\sum_{i=1}^{n}\delta_{\theta_{i}}\right)^{\otimes n}$ is i.i.d.
Corollary 1.4.
The following inequalities hold for the mutual information $I(\vartheta;X)$ and $I(\theta;X)$:
  1. 1.
    For deterministic $\theta=(\theta_{1},\dots,\theta_{n})$,
    00 0 $\leq nI(\vartheta_{1};X_{1})-I(\vartheta;X)$
    $\leq\min\left\{10\sum_{\ell=2}^{n}I_{\chi^{2}}(\vartheta_{1};X_{1})^{\ell},I_{\chi^{2}}(\vartheta_{1};X_{1})(1+\log\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))\right\}.$ (7)
  2. 2.
    For possibly random $\theta=(\theta_{1},\dots,\theta_{n})$,
    $I(\theta;X)\leq\inf_{\varepsilon>0}\left(n\varepsilon^{2}+\log N_{\mathrm{KL}}(\mathrm{conv}({\mathcal{P}}),\varepsilon)\right)+{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})(1+\log\mathsf{\Delta}_{H^{2}}({\mathcal{P}})),$ (8)
    where $\mathrm{conv}({\mathcal{P}})$ denotes the convex hull of ${\mathcal{P}}$, and
    $N_{\mathrm{KL}}({\mathcal{P}},\varepsilon):=\min\left\{m:\inf_{{\mathcal{P}}_{0}\subseteq{\mathcal{P}}:|{\mathcal{P}}_{0}|=m}\sup_{P\in{\mathcal{P}}}\min_{P_{0}\in{\mathcal{P}}_{0}}{\sf KL}(P\|P_{0})\leq\varepsilon^{2}\right\}$
    is the $\varepsilon$-covering number of the distribution class ${\mathcal{P}}$ under the KL divergence.
In words, (1) implies that the difference in the mutual information $I(\vartheta;X)$ under a permutation prior and an i.i.d. prior on $\vartheta$ is typically independent of $n$, i.e., these priors are “equally informative” for the output $X$ in the asymptotic sense. This implication turns out to find statistical applications in problems with permutation priors; we refer to Section 5.2 for such an example in the Gaussian sequence model. The inequality (8) is an upper bound on the capacity of the noisy permutation channel, where the quantity $\inf_{\varepsilon>0}\left(n\varepsilon^{2}+\log N_{\mathrm{KL}}(\mathrm{conv}({\mathcal{P}}),\varepsilon)\right)$ is an entropic upper bound on the minimax redundancy of $\mathrm{conv}({\mathcal{P}})^{\otimes n}$ [HO97, YB99], which, by the capacity-redundancy theorem [Kem74], is the channel capacity of an i.i.d. channel $X\sim Q_{\theta}^{\otimes n}$ with $Q_{\theta}\in\mathrm{conv}({\mathcal{P}})$. We note that this entropic upper bound is tight when the KL metric entropy $\log N_{\mathrm{KL}}$ is of the same order as the Hellinger metric entropy [PW24, Theorem 32.5], which holds for Gaussian or Gaussian mixture models [Ney22, Mou23, JPW23]. The inequality (8) claims that the capacity of a noisy permutation channel is essentially upper bounded by its counterpart when the conditional distribution $P_{X|\theta}$ is i.i.d., up to an $O_{\mathcal{P}}(1)$ additive factor. This generalizes the capacity upper bound proved in [TP23] for the finite-dimensional case $\dim({\mathcal{P}})=d$, where $\inf_{\varepsilon>0}\left(n\varepsilon^{2}+\log N_{\mathrm{KL}}(\mathrm{conv}({\mathcal{P}}),\varepsilon)\right)\sim\frac{d}{2}\log n$.
We mention two additional applications of Theorem 1.2 and its related proof techniques. The first application is a “noisy” finite de Finetti theorem concerning the statistical distance between the output of a kernel applied to a finite exchangeable sequence and a proper mixture of i.i.d. sequences. Let $P_{X^{n}}$ be an exchangeable distribution over ${\mathcal{X}}^{n}$, and ${\mathsf{K}}$ be a probability kernel from ${\mathcal{X}}$ to ${\mathcal{Y}}$ represented by a class of conditional distributions ${\mathcal{P}}:=\{{\mathsf{K}}(\cdot|x):x\in{\mathcal{X}}\}$. Passing $X^{n}$ through the kernel ${\mathsf{K}}^{\otimes n}$ leads to an exchangeable distribution $P_{Y^{n}}$, and the following theorem establishes its statistical distance from a mixture of i.i.d. distributions.
@@
Let $1\leq k\leq n$ and let $P_{Y^{k}}$ and ${\mathcal{P}}$ be defined as above. There exists a mixture of i.i.d. distributions $Q_{Y^{k}}=\mathbb{E}_{\theta}[Q_{\theta}^{\otimes k}]$ such that
$\chi^{2}\left(P_{Y^{k}}\|Q_{Y^{k}}\right)\leq\frac{k(k-1)}{n(n-1)}\cdot\begin{cases}10\sum_{\ell=2}^{n}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})^{\ell}&\text{if }{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})<\infty,\\ (e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}-1&\text{if }{\mathsf{C}}_{\chi^{2}}({\mathcal{P}}),\mathsf{\Delta}_{H^{2}}({\mathcal{P}})<\infty.\end{cases}$
In the absence of a smoothing kernel, some well-known results due to Stam [Sta78], Diaconis–Freedman [DF80], and several recent works [GK21, JGK24] yield bounds of the form ${\sf KL}(P_{X^{k}}\|Q_{X^{k}})=O(\frac{k^{2}}{n})$, which can be improved to $O(\frac{k^{2}}{n(n+1-k)})$ when $|{\mathcal{X}}|=O(1)$. In particular, $P_{X^{k}}$ and $Q_{X^{k}}$ are, in general, statistically indistinguishable when $k=o(\sqrt{n})$, and that this requirement can be improved to $k=o(n)$ when ${\mathcal{X}}$ is small. [^1] Note that if no restriction is made on the size of ${\mathcal{X}}$, then the birthday paradox shows that the requirement $k=o(\sqrt{n})$ is tight. In comparison, Theorem 1.5 implies that as long as the exchangeable distribution can be written as a convolution with a kernel ${\mathsf{K}}$ satisfying $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})<\infty$ (such as a Gaussian), the mixtures are statistically indistinguishable for any $k=o(n)$, without any restriction on $|{\mathcal{X}}|$. This observation suggests that the addition of noise in the form of the kernel ${\mathsf{K}}$ has a similar qualitative effect as a bound on the size of ${\mathcal{X}}$; we leave investigation of this interesting connection to future work. A similar upper bound on the KL divergence for the noisy setting also appeared in [TP23, Theorem 2], with a linear dependence on both $\frac{k}{n}$ and the cardinality of ${\mathcal{P}}$; we will provide a detailed comparison in Section 1.3.
The second application concerns the statistical distance between two different permutation mixtures. Given $P_{1},\cdots,P_{n}\in{\mathcal{P}}$, let $Q_{1}\in{\mathcal{P}}$ be an arbitrary distribution, and $Q_{i}=P_{i}$ for all $i\geq 2$. The permutation mixtures $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ are then defined as
@@
$(X_{1},\cdots,X_{n})$ $\sim\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\otimes_{i=1}^{n}Q_{\pi(i)}\right]\text{ under }\mathbb{Q}_{n}.$
In other words, the components of $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ only differ in one coordinate. The following theorem establishes an upper bound on the squared total variation (TV) distance.
@@
${\sf TV}(\mathbb{P}_{n},\mathbb{Q}_{n})^{2}\leq\frac{1}{4}\int\frac{(\mathrm{d}\mathbb{P}_{n}-\mathrm{d}\mathbb{Q}_{n})^{2}}{\mathrm{d}\overline{P}^{\otimes n}}\leq\frac{3{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{3{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}}{n},$

### v3

@@ Abstract
We prove bounds on statistical distances between high-dimensional exchangeable mixture distributions (which we call permutation mixtures) and their i.i.d. counterparts. Our results are based on a novel method for controlling $\chi^{2}$ divergences between exchangeable mixtures, which is tighter than the existing methods of moments or cumulants. At a technical level, a key innovation in our proofs is a new Maclaurin-type inequality for elementary symmetric polynomials of variables that sum to zero and an upper bound on permanents of doubly-stochastic positive semidefinite matrices. We obtain as a corollary a new de Finetti-style theorem (in the language of Diaconis and Freedman, 1987), as well as several new statistical results, including a differential privacy guarantee for the “shuffled privacy model” with Gaussian noise and improved generic consistency guarantees for empirical Bayes procedures in compound decision problems.
@@ 1 Introduction
  • Under $\mathbb{P}_{n}$, the observation $X$ is of the form $\vartheta+Z$, where $\vartheta$ is uniformly distributed on the subset of balanced vectors in $\{\pm\mu\}^{n}$ (i.e., vectors with exactly $n/2$ entries equal to $\mu$ and $n/2$ equal to $\mu$) and $Z\sim{\mathcal{N}}(0,I_{n})$ is independent Gaussian noise;
  • Under $\mathbb{Q}_{n}$, the observation $X$ is of the form $\vartheta+Z$, where $\vartheta$ is uniformly distributed on $\{\pm\mu\}^{n}$ and $Z\sim{\mathcal{N}}(0,I_{n})$ is independent Gaussian noise. Note that the entries of $X$ under $\mathbb{Q}_{n}$ are themselves i.i.d., and that the marginal distribution of each coordinate of $X$ under $\mathbb{Q}_{n}$ and $\mathbb{P}_{n}$ is the same.
@@
The law of large numbers implies that a random vector $\vartheta$ with i.i.d. $\{\pm\mu\}$ entries will be nearly balanced, and, if $\mu\lesssim 1$, the Gaussian noise will make it difficult to detect whether $\vartheta$ is perfectly balanced or only nearly so. It is therefore natural to conjecture that $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ will not be distinguishable from each other as long as $\mu$ is not too large. Surprisingly, however, as we show in Section˜3, existing approaches to bounding statistical distances between mixtures fail to establish this fact. The goal of this paper is to develop new strategies for analyzing such high-dimensional mixtures, which achieve tight bounds.
@@
Strikingly, (1) implies that the $\chi^{2}$ divergence between these two high-dimensional models is bounded independent of the dimension nn—in particular, the sequence $\mathbb{P}_{n}$ is contiguous to $\mathbb{Q}_{n}$ in the sense of Le Cam [LC60], so that asymptotic properties of the mixture $\mathbb{P}_{n}$ can be analyzed under the simpler product measure $\mathbb{Q}_{n}$. The bound in (1) also reveals other features: since $\chi^{2}({\mathcal{N}}(+\mu,1)\|{\mathcal{N}}(-\mu,1))=e^{4\mu^{2}}-1$, the $\chi^{2}$ divergence between the two models is polynomially smaller than the “single-letter” $\chi^{2}$ divergence that would arise from comparing the two univariate marginal distributions that appear in the definition of $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$. As we shall show, up to constants, the bounds in (1) are tight.
@@
Permutation mixtures arise, explicitly or implicitly, in a number of statistical settings in which either the data or parameter space enjoy permutation symmetry. However, the lack of general techniques for analyzing such mixtures has meant that researchers have had to turn to ad hoc arguments, many of which lead to overly complicated proofs or require unnecessary assumptions. The tools we develop here allow many of these results to be strengthened and simplified. In Section˜2, we illustrate this phenomenon with vignettes from high-dimensional statistics, differential privacy, and empirical Bayes estimation, showing how our bounds improve and clarify existing statistical results.
More generally, the study of permutation mixtures is motivated by important theoretical questions in a number of areas:
de Finetti-style theorems. The classical de Finetti theorem [DF29] asserts that an infinite exchangeable sequence of binary random variables is a mixture of i.i.d. Bernoulli sequences. Since the work of Diaconis [Dia77], there has been interest in obtaining quantitative forms of this theorem for finite exchangeable sequences [Sta78, DF80, GK21], and, more broadly, to obtain “de Finetti-style theorems” in general settings [DF87]. Our main results, which establish that the exchangeable measure $\mathbb{P}_{n}$ is close to the product measure $\mathbb{Q}_{n}$, are theorems of this type. As Theorem˜1.4 shows, our bounds yield new approximation results for general exchangeable sequences.
@@
Information geometry of high-dimensional mixture models. High-dimensional mixtures—such as the measures $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ defined above—are ubiquitous models in theoretical statistics (see, e.g., [LNS99, CL11, IS12, WY16, JHW18]). The analysis of such measures is routine in the univariate case (or, more generally, for product measures), but obtaining sharp bounds is difficult for high-dimensional mixtures whose coordinates are not independent. The setting we study is one in which we can rigorously compare $\mathbb{P}_{n}$ to a simpler, independent counterpart $\mathbb{Q}_{n}$. This comparison may be directly relevant in analyzing mixtures with permutation invariance. Moreover, as we highlight in Section˜1.2, our proofs are based on new techniques for bounding divergences between high-dimensional mixtures, which may be of broader applicability.
@@
$\chi^{2}(P\|Q)=\int\frac{(\mathrm{d}P-\mathrm{d}Q)^{2}}{\mathrm{d}Q},\qquad H^{2}(P,Q)=\int\left(\sqrt{\mathrm{d}P}-\sqrt{\mathrm{d}Q}\right)^{2}.$
@@
When the $\chi^{2}$ divergence is replaced by the Kullback–Leibler (KL) divergence in Definition˜1.1, the $\chi^{2}$ channel capacity coincides with the usual notion of channel capacity in information theory. We have ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})$ by convexity, and $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})\leq{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})+1$ by [PW25, Eqn. (7.33)]. In addition, $H^{2}(P_{1},P_{2})=2$ if and only if $P_{1}$ and $P_{2}$ are mutually singular, in which case it is easy to see that $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ must grow with $n$ in the worst case (cf. Lemma˜6.1). Therefore, the quantity $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})$ measures the extent of singularity of $\mathcal{P}$}. Finally, we note that all quantities (whenever finite) depend only on $\mathcal{P}$} and not on the dimension $n$ of $\mathbb{P}_{n}$ or $\mathbb{Q}_{n}$. We present our main theorem below, which upper bounds the $\chi^{2}$ divergence using only the above quantities.
@@
Theorem˜1.2 shows that if ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq 1-\delta$ for some constant $\delta>0$, then the $\chi^{2}$ divergence is upper bounded by a dimension-independent constant $O_{\delta}({\mathsf{C}}_{\chi^{2}}({\mathcal{P}})^{2})$. In addition, even if ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\geq 1$ but $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})<\infty$, the $\chi^{2}$ divergence is still upper bounded by a large albeit dimension-independent constant. This shows that the contiguity relation $\{\mathbb{P}_{n}\}\triangleleft\{\mathbb{Q}_{n}\}$ is not unique to the Gaussian family but holds for a large class of probability families $\mathcal{P}$}. Specializing $\mathcal{P}$} to specific classes of distributions leads to the following corollary.
@@
The proof of Corollary˜1.3 follows from Theorem˜1.2 and the computations of $({\mathsf{C}}_{\chi^{2}}({\mathcal{P}}),\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))$, which we defer to Section˜C.2. For the example presented at the beginning of the introduction, a weaker upper bound $\exp(O(\mu^{2}))$ of (1) follows from point 2 of Corollary˜1.3; we refer to Section˜4.1 for the proof of the stronger form (1) (via the inequality ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq 1-\exp(-\mu^{2})$ in (2)). We will also discuss the tightness of these upper bounds in Section˜6.1; typically, the quadratic dependence in (4) is tight when ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ is small, and both the base and exponent of (5) are not improvable in general. However, for specific $\mathcal{P}$} there could still be gaps, and we refer the discussions to Section˜6.1.
Theorem˜1.2 admits several generalizations to related settings.
A “noisy” finite de Finetti theorem.
The same proof technique of Theorem˜1.2 also yields a version of the de Finetti theorem for finite exchangeable sequences with added noise. Given a permutation mixture $\mathbb{P}_{n}$ and its i.i.d. counterpart $\mathbb{Q}_{n}$, we denote by $\mathbb{P}_{k,n}$ and $\mathbb{Q}_{k,n}$ the marginal distribution of the first $k$ coordinates:
$\mathbb{P}_{k,n}$ $=\mathrm{Law}(X_{1},\dots,X_{k})\quad\quad\text{for $(X_{1},\dots,X_{n})\sim\mathbb{P}_{n}$}$
$\mathbb{Q}_{k,n}$ $=\mathrm{Law}(X_{1},\dots,X_{k})\quad\quad\text{for $(X_{1},\dots,X_{n})\sim\mathbb{Q}_{n}$}\,.$
Of course, $\mathbb{Q}_{k,n}$ is still a product distribution. The following theorem establishes bounds on the statistical distance between $\mathbb{P}_{k,n}$ and this i.i.d. counterpart.
@@
Let $1\leq k\leq n$. Then
$\chi^{2}\left(\mathbb{P}_{k,n}\|\mathbb{Q}_{k,n}\right)\leq\frac{k^{2}}{n^{2}}\cdot\begin{cases}20{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})^{2}&\text{if }{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq\frac{n}{2k},\\ (e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}-1&\text{if }{\mathsf{C}}_{\chi^{2}}({\mathcal{P}}),\mathsf{\Delta}_{H^{2}}({\mathcal{P}})<\infty.\end{cases}$
By a standard convexity argument, the same bound holds when comparing the marginal distribution of an arbitrary finite exchangeable sequence (which can be written as a mixture of permutation mixtures) to a mixture of product distributions.
In the absence of upper bounds on ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ and $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})$, some well-known results due to Stam [Sta78], Diaconis–Freedman [DF80], and several recent works [GK21, JGI25] yield bounds of the form ${\sf KL}(\mathbb{P}_{k,n}\|\mathbb{Q}_{k,n})=O(\frac{k^{2}}{n})$, and $O(\frac{k^{2}}{n(n+1-k)})$ when $|{\mathcal{P}}|=O(1)$. In particular, $\mathbb{P}_{k,n}$ and $\mathbb{Q}_{k,n}$ are, in general, statistically indistinguishable when $k=o(\sqrt{n})$, and this requirement can be improved to $k=o(n)$ when $\mathcal{P}$} is small. [^1] Note that if no restriction is made on the size of $\mathcal{P}$} or on $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})$, then the birthday paradox shows that the requirement $k=o(\sqrt{n})$ is tight. Theorem˜1.4 generalizes these results from two directions. First, if $|{\mathcal{P}}|=O(1)$, we have ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq|{\mathcal{P}}|-1$ (corollary of Lemma˜C.1) even when the elements of $\mathcal{P}$} are mutually singular. Therefore, our first upper bound gives $\chi^{2}(\mathbb{P}_{k,n}\|\mathbb{Q}_{k,n})=O(\frac{k^{2}}{n^{2}})$ for $k\leq cn$, generalizing Stam’s results [Sta78] to the $\chi^{2}$ divergence, which cannot be analyzed using his techniques. Second, the same $O(\frac{k^{2}}{n^{2}})$ scaling also holds when $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})<\infty$. This corresponds to the setting in which the distributions in $\mathcal{P}$} are “noisy” enough that the pairwise squared Hellinger distances are bounded away from $2$. In this case, the mixtures are statistically indistinguishable for any $k=o(n)$, without any restriction on $|{\mathcal{P}}|$. This observation suggests that the addition of noise in the form of a bound on $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})$ has a similar qualitative effect as a bound on the size of $\mathcal{P}$}.
Neighboring permutation mixtures.
The next result concerns the statistical distance between two “neighboring” permutation mixtures. Given $P_{1},\cdots,P_{n}\in{\mathcal{P}}$, let $P^{\prime}_{1}\in{\mathcal{P}}$ be arbitrary and set $P^{\prime}_{i}=P_{i}$ for all $i\geq 2$. The permutation mixtures $\mathbb{P}_{n}$ and $\mathbb{P}^{\prime}_{n}$ are then defined as
@@
$(X_{1},\cdots,X_{n})$ $\sim\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\otimes_{i=1}^{n}P^{\prime}_{\pi(i)}\right]\text{ under }\mathbb{P}^{\prime}_{n}.$
In other words, the components of $\mathbb{P}_{n}$ and $\mathbb{P}^{\prime}_{n}$ only differ in one coordinate. The following theorem establishes an upper bound on the squared total variation (TV) distance.
@@
${\sf TV}(\mathbb{P}_{n},\mathbb{P}^{\prime}_{n})^{2}\leq\frac{1}{4}\int\frac{(\mathrm{d}\mathbb{P}_{n}-\mathrm{d}\mathbb{P}^{\prime}_{n})^{2}}{\mathrm{d}\overline{P}^{\otimes n}}\leq\frac{3{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{3{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}}{n},$